I need new and unique answers, please. (Use your own words, don't copy and paste), Please Use your keyboard (Don't use handwriting) Thank you..

Q1-

Abdulaziz Co. purchased a machine in 2013 for 50,000 that has a useful life of 5 years with a salvage value of 5,000.

Calculate the depreciation expense, accumulated depreciation, book value throughout its useful life using:

1- Straight-line Method.

2- Units of Production Method if the machine produces 100,000 units.

Here is a table of units produced each year:

First

Second

Third

Fourth

Fifth

23,000

25,000

-

30,000

22,000

3- Double Declining Balance Method (2 Marks).

Q2- On June 1, 2019, ABC Company signed a $25,000, 120-day, 6% note payable to cover a past due account payable.
a. What is the total amount of interest to be paid on this note?
b. Prepare ABC Company's general journal entry to record the issuance of the note payable
c. Prepare ABC Company's general journal entry to record the payment of the note on
September 29, 2019 (2 marks).

Q3. What are the characteristics of corporations (1 mark).
Answer #1

Q1.
Straight line method
Depreciation expense = (Original Value - Salvage Value) / Useful Life
= ($50000-5000)/5 = $9000 per year

Year Beginning Book Value Depreciation Accumulated Depreciation Ending Book Value
2013 $            50,000 $           9,000 $            9,000 $      41,000
2014 $            41,000 $           9,000 $            9,000 $      32,000
2015 $            32,000 $           9,000 $            9,000 $      23,000
2016 $            23,000 $           9,000 $            9,000 $      14,000
2017 $            14,000 $           9,000 $            9,000 $        5,000

Units of Production Method
Depreciation expense = (Original Value - Salvage Value) / Total Capacity
= ($50000-5000)/100000 = $0.45 per unit

Year Beginning Book Value Depreciation Accumulated Depreciation Ending Book Value
2013 $            50,000 $         10,350 $          10,350 $      39,650
2014 $            39,650 $         11,250 $          11,250 $      28,400
2015 $            28,400 $                  -   $                  -   $      28,400
2016 $            28,400 $         13,500 $          13,500 $      14,900
2017 $            14,900 $           9,900 $            9,900 $        5,000

Double Declining Balance
Depreciation = Beginning book Value x 2 times straight line rate
Straight line rate = 1/5 x 100 = 20%

Year Beginning Book Value Depreciation Accumulated Depreciation Ending Book Value
2013 $            50,000 $         20,000 $          20,000 $      30,000
2014 $            30,000 $         12,000 $          12,000 $      18,000
2015 $            18,000 $           7,200 $            7,200 $      10,800
2016 $            10,800 $           4,320 $            4,320 $        6,480
2017 $              6,480 $           1,480 $            1,480 $        5,000


Depreciation for Year 2017 is lower than 40% of Beginning Book Value as Ending Book Value cannot be lower than Salvage Value.

Note : As per HOMEWORKLIB RULES we are supposed to do one question completely, so please post other questions separately